# Strong Balmer break objects at uncovered with JWST

### v1

@@ 3.3 Balmer break galaxies at are more frequent than previously thought
Out of the 101 sources for which secure spectroscopic redshift is available, we found spectra for 62 objects in DJA. Out of these 62, 5 () sources show a BB with a flat red continuum like in a mini-quenched galaxy, 5-6 () sources show a Balmer break with emission lines and one () is an LRD-like candidate with a possible break. That means that of the sources in our sample show a BB or the presence of an evolved stellar population.
@@
Alternatively, inspection of the spectra of all objects with photometric BB (16 sources) confirms the presence of a BB in 6 of them ( %), shows emission lines in 6 of them (37 %), and the remaining spectra are of insufficient quality. This translates to a failure rate of 37% (maximum 63%) to find true BB candidates from our photometric method.
@@
If we consider the full photometric sample (with secure spectroscopic redshifts), 27 sources show BB above 1.2 out of 101 (%), comparable to the inference from the spectra. If one assumes the above-mentioned failure rate for the whole sample, % will have a true BB, which is larger than the previously confirmed BB objects but comparable to the fraction of āsmoulderingā (i.e.Ā in with suppressed star-formation) galaxies around at determined by Trussler etĀ al. (2023). In any case, with a total sample of 10-11 spectroscopically-confirmed BB objects at , our work essentially triples the number of such objects.
@@ 4 Conclusions
To progress towards establishing the occurrence and properties of galaxies showing Balmer breaks (BBs) at , we have selected all galaxies with secure spectroscopic redshifts from the main extragalactic surveys (CEERS, JADES, PRIMER, FRESCO) observed with JWST for which we measure the strength of the Balmer break from photometry. We have validated our method by comparison with spectroscopic measurements of the BB from the stacked spectra of Roberts-Borsani etĀ al. (2024), and we have inspected the individual spectra to validate our measurements and assess the nature of the galaxies.
@@
Our work has revealed three new galaxies from with strong BBs and no signs of emission lines, which have most likely stopped star formation recently (so-called mini-quenched objects). We have also found 5-6 galaxies showing the co-existence of strong BBs and emission lines (H, [OĀ iii]), which we interpret as signs of renewed star formation after a previously interrupted (quenched) phase. This approximately triples the number of spectroscopically-confirmed galaxies at showing clear signs of quenched or interrupted star-formation, with respect to previous studies (Looser etĀ al. 2024; Vikaeus etĀ al. 2024; Wang etĀ al. 2024b; Witten etĀ al. 2024).
@@
Overall, from our sample of 118 galaxies, 101 yielded BB measurements, with approximately 48% showing a net BB, and % a āstrongā BB. From a detailed inspection of the available spectra (for a subsample of 62 objects), we estimate that % show a clear BB, out of which approximately half show no signs of ongoing star-formation, and the rest show BB plus emission lines. The remaining objects with apparent strong BBs resemble LRDs or cannot well be classified. Our photometric BB measurements show that objects with strong BBs can be found over a wide range of magnitudes and stellar masses. From our analysis, we conclude that the fraction of mini-quenched or reactivated (bursty) galaxies at is of the order of %, at least in the sample studied here. Future work will be needed to understand the possible selection biases, establish how these trends evolved towards lower redshift, quantify the timescales involved in the observed star-formation and intermittent phases, and work out more general implications.

### v2

@@ 3.3 Balmer break galaxies at are more frequent than previously thought
Out of the 101 sources for which secure spectroscopic redshift is available, we found spectra for 62 objects in DJA. Out of these 62, 5 () sources show a BB with a flat red continuum like in a mini-quenched galaxy, 5-7 () sources show a Balmer break with emission lines and one () is an LRD-like candidate with a possible break. That means that of the sources in our sample show a BB or the presence of an evolved stellar population.
@@
Alternatively, an inspection of the spectra of all objects with photometric BB (16 sources) yields the following: a BB is clearly detected and thus confirmed in 6 of them ( %), 6 others show emission lines and no clear break (37 %), and the remaining spectra are of insufficient quality. This translates to a failure rate of 37% (maximum 63%) to find true BB candidates from our photometric method.
@@
If we consider the full photometric sample (with secure spectroscopic redshifts), 27 sources show BB above 1.2 out of 101 (%), comparable to or slightly higher than the inference from the spectra. If one assumes the above-mentioned failure rate (and range thereof) for the whole sample, % will have a true BB, which is larger than the previously confirmed BB objects but comparable to the fraction of “smouldering” (i.e. in with suppressed star-formation) galaxies around at determined by Trussler et al. (2024a). In any case, with a total sample of 10-11 spectroscopically confirmed BB objects at , our work essentially triples the number of such objects.
Previously, JWST spectroscopic observations revealed 4 objects with significant Balmer breaks at (Looser et al. 2024; Vikaeus et al. 2024; Wang et al. 2024b; Witten et al. 2024) and a galaxy with a BB 1 (Boyett et al. 2024), as also shown in Figs. 1 and 4 333One additional galaxy with spectroscopic redshift and BB is reported in the work of Trussler et al. (2024a) mentioned above.. These studies reported individual objects selected by different criteria, and which do not result from a systematic search for strong BB objects. Their BB measurements from spectroscopy agree well (within 1-2 ) with ours from photometry, as seen from Fig. 1, and the spectroscopic measurements are not affected by possible emission lines. Our sample not only adds new spectroscopically-confirmed BB objects, but it also extends somewhat the magnitude range over which such objects are found, as shown in 4.
3.4 Origin of the observed Balmer breaks
None of the objects reported here shows a Balmer break exceeding any fundamental limit. In particular, the observed values of BB are within the range of predicted for normal stellar populations and can be reached with ages less than the Hubble time, as shown earlier by Wang et al. (2024b) for the strongest BB object studied here. The BB strengths measured in this study are in agreement with the theoretical predictions of Binggeli et al. (2019) and Wilkins et al. (2024), where some simulations predict BB strength up to 1.5 - 2.0 in redshift with extinction. Smaller values are predicted by Steinhardt et al. (2024) who use, however, D4000, which shows a small dynamical range due to chosen wavelength windows. In short, our strong BB candidates do not break standard cosmology, but they might indicate the need for new star formation models or obscured star formation (see Fig. 5 in Steinhardt et al. 2024).
Most likely, the sample presented here represents the bursty nature of galaxies of galaxies. Indeed, many theoretical and observational studies have already proposed that star formation in the early Universe might be bursty and/or stochastic (e.g. Caputi et al. 2017; Faisst et al. 2019; Rinaldi et al. 2022; Endsley et al. 2024; Looser et al. 2023; Sun et al. 2023; Looser et al. 2024; Witten et al. 2024). These starburst phases are accompanied by temporary or permanent phases of quiescence, presumably due to various feedback mechanisms arising after the burst phase, which then allows the older stellar population (and the BB) to become visible in the integrated spectra. The co-existence of BBs with emission lines probably indicates the restart of star formation after a temporary quiescence (Witten et al. 2024). In summary, the observed range of BB strengths and the existence of BB galaxies with emission lines indicate the galaxies at are transiting between on and off phases of star formation (cf. Witten et al. 2024; Looser et al. 2024; Trussler et al. 2024a). The presence of these galaxies provides constraints on the star-formation histories in the early Universe and their detailed modeling will help us to constrain the onset of galaxy formation.
A more detailed study of the evolution of the BB in large galaxy samples and over a wider redshift range will be presented in a follow-up study (Kuruvanthodi et al., in prep.).
@@ 4 Conclusions
To progress towards establishing the occurrence and properties of galaxies showing Balmer breaks (BBs) at , we have selected all galaxies with secure spectroscopic redshifts from the main extragalactic surveys (CEERS, JADES, PRIMER, FRESCO) observed with JWST for which measure the strength of the Balmer break from photometry. We have validated our method by comparison with spectroscopic measurements of the BB from the stacked spectra of Roberts-Borsani et al. (2024), and we have inspected the individual spectra to validate our measurements and assess the nature of the galaxies.
@@
Our work has revealed three new galaxies from with strong BBs and no signs of emission lines, which have most likely stopped star formation recently (so-called mini-quenched objects). We have also found 5-7 galaxies showing the co-existence of strong BBs and emission lines (H, [O iii]), which we interpret as signs of renewed star formation after a previously interrupted (quenched) phase. This approximately triples the number of spectroscopically-confirmed galaxies at showing clear signs of quenched or interrupted star formation, with respect to previous studies (Looser et al. 2024; Vikaeus et al. 2024; Wang et al. 2024b; Witten et al. 2024).
@@
Overall, from our sample of 118 galaxies, 101 yielded BB measurements, with approximately 48% showing a net BB, and % a “strong” BB. From a detailed inspection of the available spectra (for a subsample of 62 objects), we estimate that % show a clear BB, out of which approximately half show no signs of ongoing star formation, and the rest show BB plus emission lines. The remaining objects with apparent strong BBs resemble LRDs or cannot well be classified. Our photometric BB measurements show that objects with strong BBs can be found over a wide range of magnitudes and stellar masses. From our analysis, we conclude that the fraction of mini-quenched or reactivated (bursty) galaxies at is of the order of %, at least in the sample studied here.
Our sample, showing a range of BB strength and line properties, highlights the bursty and stochastic nature of star formation in the early Universe and the existence of significant galaxies in a temporary phase of quiescence in the high-z Universe. Future work will be needed to understand the possible selection biases, establish how these trends evolved towards lower redshift, quantify the timescales involved in the observed star formation and intermittent phases, and work out more general implications.
Acknowledgements.
We thank the anonymous referee for her/his insightful comments and suggestions. The data products presented herein were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. We thank Gabe Brammer for making this tool available. We thank Prof. Corinne Charbonnel for her useful comments and suggestions. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349References
- Binggeli et al. (2019) Binggeli, C., Zackrisson, E., Ma, X., et al. 2019, MNRAS, 489, 3827
- Boyett et al. (2024) Boyett, K., Trenti, M., Leethochawalit, N., et al. 2024, Nature Astronomy, 8, 657
- Brammer (2023) Brammer, G. 2023, msaexp: NIRSpec analyis tools
- Caputi et al. (2017) Caputi, K. I., Deshmukh, S., Ashby, M. L. N., et al. 2017, The Astrophysical Journal, 849, 45
- Eisenstein et al. (2023) Eisenstein, D. J., Willott, C., Alberts, S., et al. 2023, arXiv e-prints, arXiv:2306.02465
- Endsley et al. (2024) Endsley, R., Stark, D. P., Whitler, L., et al. 2024, MNRAS, 533, 1111
- Eyles et al. (2007) Eyles, L. P., Bunker, A. J., Ellis, R. S., et al. 2007, MNRAS, 374, 910
- Faisst et al. (2019) Faisst, A. L., Capak, P. L., Emami, N., Tacchella, S., & Larson, K. L. 2019, The Astrophysical Journal, 884, 133
- Finkelstein et al. (2023) Finkelstein, S. L., Bagley, M. B., Ferguson, H. C., et al. 2023, ApJL, 946, L13
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
- Hashimoto et al. (2018) Hashimoto, T., Laporte, N., Mawatari, K., et al. 2018, Nature, 557, 392
- Heintz et al. (2024) Heintz, K. E., Brammer, G. B., Watson, D., et al. 2024, arXiv e-prints, arXiv:2404.02211
- Kocevski et al. (2024) Kocevski, D. D., Finkelstein, S. L., Barro, G., et al. 2024, arXiv e-prints, arXiv:2404.03576
- Labbe et al. (2023) Labbe, I., Greene, J. E., Bezanson, R., et al. 2023, arXiv e-prints, arXiv:2306.07320
- Laporte et al. (2017) Laporte, N., Ellis, R. S., Boone, F., et al. 2017, ApJL, 837, L21
- Laporte et al. (2023) Laporte, N., Ellis, R. S., Witten, C. E. C., & Roberts-Borsani, G. 2023, MNRAS, 523, 3018
- Looser et al. (2023) Looser, T. J., D’Eugenio, F., Maiolino, R., et al. 2023, arXiv e-prints, arXiv:2306.02470
- Looser et al. (2024) Looser, T. J., D’Eugenio, F., Maiolino, R., et al. 2024, Nature, 629, 53
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- Mawatari et al. (2020) Mawatari, K., Inoue, A. K., Hashimoto, T., et al. 2020, The Astrophysical Journal, 889, 137
- Oesch et al. (2023) Oesch, P. A., Brammer, G., Naidu, R. P., et al. 2023, MNRAS, 525, 2864
- Rinaldi et al. (2022) Rinaldi, P., Caputi, K. I., van Mierlo, S. E., et al. 2022, The Astrophysical Journal, 930, 128
- Roberts-Borsani et al. (2024) Roberts-Borsani, G., Treu, T., Shapley, A., et al. 2024, arXiv e-prints, arXiv:2403.07103
- Roberts-Borsani et al. (2020) Roberts-Borsani, G. W., Ellis, R. S., & Laporte, N. 2020, Monthly Notices of the Royal Astronomical Society, 497, 3440
- Schaerer & de Barros (2009) Schaerer, D. & de Barros, S. 2009, A&A, 502, 423
- Steinhardt et al. (2024) Steinhardt, C. L., Sneppen, A., Clausen, T., et al. 2024, The Astrophysical Journal, 967, 172
- Strait et al. (2023) Strait, V., Brammer, G., Muzzin, A., et al. 2023, ApJL, 949, L23
- Sun et al. (2023) Sun, G., Faucher-Giguère, C.-A., Hayward, C. C., & Shen, X. 2023, MNRAS, 526, 2665
- Trussler et al. (2024a) Trussler, J., Conselice, C., Adams, N., et al. 2024a, arXiv e-prints, arXiv:2404.07163
- Trussler et al. (2024b) Trussler, J. A. A., Conselice, C. J., Adams, N., et al. 2024b, MNRAS, 527, 11627
- Vikaeus et al. (2024) Vikaeus, A., Zackrisson, E., Wilkins, S., et al. 2024, MNRAS, 529, 1299
- Wang et al. (2024a) Wang, B., de Graaff, A., Davies, R. L., et al. 2024a, arXiv e-prints, arXiv:2403.02304
- Wang et al. (2024b) Wang, B., Leja, J., de Graaff, A., et al. 2024b, ApJL, 969, L13
- Weibel et al. (2024) Weibel, A., Oesch, P. A., Barrufet, L., et al. 2024, MNRAS, 533, 1808
- Wilkins et al. (2024) Wilkins, S. M., Lovell, C. C., Irodotou, D., et al. 2024, MNRAS, 527, 7965
- Witten et al. (2024) Witten, C., McClymont, W., Laporte, N., et al. 2024, arXiv e-prints, arXiv:2407.07937
Appendix A Spectra of other Balmer break candidates
Spectra of other BB candidates with emission lines are shown here.
Appendix B Coordinates and properties of the Balmer break sources/candidates
ID | RA | Dec | BB | |||
(deg) | (deg) | (mag) | (Å) | |||
7.26 | 53.1528407 | -27.8019455 | ¡21.55 | |||
7.48 | 189.1211092 | 62.2778222 | ¡59.36 | |||
8.53 | 189.1807727 | 62.1805329 | ¡36.52 | |||
9.56 | 214.8118531 | 52.7371100 | ¡137.80 | |||
7.48 | 214.9278186 | 52.8500015 | ||||
7.14 | 150.0990078 | 2.3436228 | ||||
7.27 | 53.1910569 | -27.7973172 | ||||
7.27 | 53.1837525 | -27.7938918 | ¡148.60 | |||
7.48 | 214.8120544 | 52.7467436 | ||||
7.65 | 214.8060747 | 52.7508628 | ||||
8.66 | 189.2718353 | 62.1951783 | ¡31.61 | |||
8.68 | 53.1760306 | -27.8173945 | ¡72.98 |